# Instability and stress fluctuations of a probe driven through a worm-like micellar fluid

### v1

@@ III Results and discussions
To gain insights into the origin of the unstable flows in the system, we study the flow-induced sample turbidity using in-situ transmission-mode imaging [Fig. S4]. In WLM systems, sample turbidity occurs due to shear-induced local density variation over a range of length scales, enabling visible light scattering[3, 40, 41]. We perform these experiments in the separate motor transducer (SMT) mode where the pin is held fixed and the bottom sample cell is rotated at a desired angular velocity with the help of the bottom motor. Figs. 2(b)-(d) show the snapshots of the sample turbidity in steady-state corresponding to fixed applied velocities of 2.2 mm/s, 9 mm/s, and 15 mm/s, respectively. For v = 2.2 mm/s, we do not see any flow-induced structure and the image shows a uniform darkness. For v = 9 mm/s, we observe the formation of a long bright extended wake that remains attached to the pin. The bright band-like structure remains stable under flow without showing any length variation. Interestingly, right after the beginning of the plateau region of the F-v plot, such a bright band starts to show continuous attachment-detachment events with the pin [Fig. 2(d), see movie 2]. The onset velocity of the pin to observe such events correlates with that of the stress fluctuation events in a one-to-one manner. This indicates that the attachment/detachment dynamics of the flow-induced structures are intricately related to the dynamics of steady-state force fluctuations. Interestingly, such sample turbidity under shear (Fig. 2b - 2d) does not depend on the polarization of light. We have tested this by measuring turbidity by placing a linear polarizer in front of the light source and an analyzer in front of the camera. We observe that the sample turbidity for a particular velocity of the pin (in the force-plateau region) remains unchanged for different combinations of the optical axes of the polarizer and the analyzer. Our observations are consistent with the sample turbidity obtained from bulk rheology measurements on the same sample, where no polarization dependence is observed [16].
@@
Next, to establish a correlation between the attachment detachment dynamics with the observed force fluctuations, we map out the velocity distribution in the system using the particle imaging velocimetry (PIV) technique. As the micellar solution is transparent, 1.5 wt% of polystyrene tracer particles of size varying between 20-50 are added to the system to generate the contrast required for PIV. We confirm that adding such a small amount of tracer particles does not modify the rheological properties of the WLM system under study [Fig. S1]. For visualization, we image the top surface of the sample with the pin held fixed and the bottom geometry rotated [Fig. S5]. The recorded images are analyzed using the PIV Lab algorithm using MATLAB software[36]. Fig. 3(a) shows the typical snapshot of the raw image of the sample-air interface seeded with tracer particles. Here we note that due to a difference between the radial distance of the two walls of the annular geometry from the axis of rotation, even without the presence of the pin the azimuthal velocity increases linearly from the inner to the outer wall. Fig. 3(b) shows the spatial distribution of velocity in the steady state for v = 2.5 mm/s. The velocity distribution in the presence of the pin (depicted as a black circle in the velocity distribution map) is almost similar to that without the pin. This indicates that for low velocities well below the onset of the plateau of the vs curve, the pin perturbs the system only locally, as also seen by a small low-velocity region around the pin [Fig. 3(b)]. When the applied velocity is increased to 12 mm/s (just below the onset of the plateau shown in Fig. 3(c)), a wake having a long, tail-like stationary structure is formed that remains connected to the probe and has much smaller velocity magnitude compared to the other portion of the sample. Interestingly, in the case of Newtonian fluids like glycerol, we observe that the presence of the pin perturbs the velocity distribution of the system only locally even if the velocity is varied over a wide range [Fig. S17]. This marks the striking difference between the flow fields around the pin for WLM system as compared to a high-viscosity Newtonian liquid. As mentioned earlier, we obtain large temporal fluctuations of F for an applied velocity that lies in the early plateau region of the curve where the features of the fluctuations are saw-tooth type. In Fig. 3(d), we show a zoomed-in version of a typical fluctuation event for v = 15 mm/s showing a slow build-up followed by a sharp drop in force. As indicated in the figure, we mark different regions in the time series as i, ii, iii, and iv. The corresponding spatial distributions of velocity are shown in Fig. 3 (panels i - iv). Just before the sharp force drop (region i), the spatial distribution of velocity shows the formation of a flow-induced long wake-like structure of a low-velocity band behind the pin (similar to the stable wake obtained for v = 12 mm/s). Such a stable wake quickly detaches from the pin and starts to move away from the pin with a much higher velocity compared to the velocity of the moving sample cell. At this point, force also starts to drop. Subsequently, the detached wake disturbs the sample around and pushes until a maximum speed of the sample is reached (iii and iv) and also reaches a minimum value. After this, the formation of a stable wake connected with the pin again takes place. The gradual increase in takes place correlating with the gradual elongation of the wake-like structure until the next force-drop event sets in. Such events continue as long as the sample cell velocity is maintained. Our results establish that the force fluctuation events and the formation and detachment of flow-induced structures are correlated in a one-to-one manner.
@@
In the increasing branch of the saw-tooth waveform for [Fig. 4(a)], we find that the length of the wake attached to the pin also increases with time till the maximum value of the force is reached (before the sharp force drop). Here, we define the length of the wake as the length of the region along the direction of azimuthal velocity over which the flow field gets altered due to the presence of the pin. Physically, this length scale signifies the region of influence of the pin for a given driving velocity and is obtained from the PIV measurement [panel i and ii in Fig. 4(a)]. We denote the maximum length of the wake formed within a saw-tooth waveform as . We further define a rise-time as the time required for to rise monotonically from a local minimum value to reach a maximum before it starts to drop abruptly as shown in Fig. 4(a). Earlier studies indicate that changing the degree of screening by varying the salt concentrations can tune the flow-concentration coupling in CTAT WLM[18]. We observe that a small variation of salt concentration (30 mM to 100 mM NaCl) does not change the nature of the curve except for slight shifting of the plateau onset [Fig. S10]. We also find a similar saw-tooth wave pattern close to the plateau onset for different salt concentrations. We construct a length scale from the force rise time: , where is the set velocity of the annular geometry at the location of the pin. We observe that (obtained from PIV measurement) and (obtained from the saw tooth behaviour of time-varying force) are very close to each other and vary in a correlated fashion as shown in Fig. 4(b) for a range of salt concentrations. Interestingly, we observe from Fig. 4(b) that the length of the wake before detachment increases with increasing salt concentration. Our results highlight that the rise of the force on the probe originates from the stretching of the micellar structure in the wake and the detachment of the structure from the pin results in the sharp drop in the force. Although close to the plateau region of the curve, the nature of most fluctuation events is quite close to the saw-tooth shape, few of the events are more complex (e.g. the first event shown in Fig. 4(a)) showing local non-monotonicity. In such cases, the wake region shows a partial detachment. We ignore such events in our correlation plot in Fig. 4(b). Interestingly, we observe that such partial detachment events become progressively more frequent as we enter deeper inside the plateau region. This observation rationalizes the complex waveform observed in the deep plateau region giving rise to more turbulent/chaotic flows. Apart from the change in salt concentration, temperature also plays a significant role in controlling the structural and mechanical properties of WLM. We find that the overall nature of the curve and the nature of fluctuations remain similar over a range of temperatures from to that we probe. We observe that for our system, the plateau onset shifts towards higher velocity with increasing temperature [Fig. S9]. However, estimating the length of the wake using PIV analysis at different temperatures remains outside the scope of the present study due to the technical limitation of controlling the sample temperature in twin-drive mode.
@@
The extensional deformation response of the system at room temperature of is probed using a custom made dripping on to substrate (DoS) rheometry technique following the standard protocol mentioned in the literature[37, 38, 39]. We observe the time evolution of a liquid bridge between a modified substrate and the needle that pinches off due to the role of strong capillary force. Fig. 6(a) shows the snapshots of the liquid bridge at different times for 2wt% CTAT + 100 mM based wormlike micellar system at room temperature of . We record the time evolution of the minimum neck diameter() of the unstable liquid bridge normalized w.r.t. the needle diameter () and extract the extensional relaxation time by fitting an exponential curve () as predicted by the elasto-capillary model[42]. Fig. 6(b) shows the temporal evolution of the normalized mid-filament diameter for 2wt% CTAT system with variation of NaCl concentrations (see Fig. S20 for snapshots in the case of 30 mM and 50 mM NaCl) and the dashed lines are the fits using elasto-capillary model. The data is reproducible over multiple set of experiments [Fig. S21]. The extensional relaxation time obtained from DoS fittings over 4 repeated measurements is plotted w.r.t. the variation of NaCl concentrations (see inset of Fig.6(b)).
@@
Interestingly, 2 wt % CTAT + 100 mM NaCl does not show shear banding (before unstable/time-varying flows the velocity profile remains linear) as also expected from the finite slope ( 0.3) of the plateau region of the flow curve [16]. At high , the system directly enters into an unstable flow regime with random time-varying changes in the flow profile. However, in our pin-driving set-up, the force-velocity curve shows a very flat plateau indicating the possibility of shear-banding. Indeed, our PIV measurement clearly shows that significant non-linearity in the flow profile develops near the plateau onset at much lower driving velocities beyond which the time-varying flow develops. Such a non-linear flow profile originates from the stable wake formation. The estimated value of Reynolds number (, where is the density, is the driving velocity, is the pin diameter, is the zero shear viscosity) at the plateau onset is much smaller than 1 () signifying that the elasticity plays a major role. Also the extensile Deborah number [30, 22] (, where is the maximum extension rate obtained from PIV measurement) at the plateau onset is 0.4 which is smaller than 1. Remarkably, shear ( ) and extensional () Weissenberg numbers ( and respectively denote the maximum shear and extension rates close to the pin) at the onset of time-varying flow in pin-driving set-up ( 3 and 3) are more than unity, but remain well below that obtained in the Taylor-Couette geometry for the same system (Wi >15). Such differences indicate that the mechanism of flow instability in the pin-driving set-up fundamentally differs from that obtained in Taylor-Couette geometry and the bulk rheological properties of the materials alone are not sufficient to describe the motion and instability of a small probe driven through the micellar fluid.
@@ IV Conclusions
In summary, we investigate the motion of a probe driven through a wormlike micellar fluid at a controlled velocity using a home-built shear cell coupled with in-situ flow visualization. For the first time, we correlate temporal stress fluctuations on the probe and the local shear and extensional deformations in a one-to-one fashion. We find that the flat force plateau obtained in the present setup differs from the finite slope of the stress plateau obtained in bulk rheology measurements for the same WLM system. We also obtain very different onset for stress fluctuations in pin-driving setup compared to that observed for a Taylor-Couette geometry. These observations highlight that bulk rheological parameters are insufficient to capture the nature of flow instability for the small probe driven through a WLM fluid. Interestingly, the flow behind a moving cylinder is geometrically quite similar to the present setup with an additional complexity of shear at the boundaries in the pin-driving geometry. There have been detailed studies of flow instability for both Newtonian, as well as, viscoelastic shear-thinning and Boger fluids [43, 44]. For viscoelastic flows, elastic wake instability and the formation of 3-D cellular structure along the length of the cylinder have been observed. In our setup, the pin completes one full rotation in less than 1 s for the pin velocity greater than 450 mm/s. For our WLM system, this sets the upper limit of the velocity as the shear relaxation time is 1 s. Even at such high velocity, the Reynolds number just crosses 1 for glycerol. Thus, we need to go to a much higher velocity to induce inertia-driven instability in glycerol. The reliability of force measurement and the high frame rate requirement will need significant optimization of the setup. Nonetheless, these directions will be very interesting to explore in the context of the motion of small probes in WLM systems in the future. Our study reveals the formation of flow-induced stable wakes just before the onset of force fluctuations. At higher driving velocities, the attachment detachment dynamics of the shear-induced micellar structures in the wake region with the probe remains strongly correlated with the stress fluctuations in the system. In the wake region, we get flow-induced turbidity which is visible even under the ambient light. This turbid band forms due to the scattering of light from the shear-induced structures in the non-linear flow regime. Interestingly, the scattered intensity is not sensitive to the polarization of light which we verify using different orientations of optical axes of a linear polarizer-analyser pair. Understanding the microscopic origin of the sample turbidity will be an interesting future research direction. Careful measurement of sample birefringence might give us important insights. In the present study, we probe the stress fluctuations under a controlled velocity. Our preliminary data on controlled stress measurements reproduce the damped harmonic oscillations and bursts observed in the ball drop experiments in the CTAT-NaCl systems[4]. Our experiments can probe burst statistics much better as the conventional ball drop experiments are limited by the finite tube length. However, we must remember that the probe is held rigidly in the current setup which does not allow any rotational or lateral motion of the probe. However, in the case of ball-drop experiments, no such restrictions are present. This can affect the dynamics. Introducing such flexibility in our set-up will be an interesting direction for future studies.
@@
See the suplementary material for information about different imaging set-ups, data analysis and additional measurements.

### v2

@@ III Results and discussions
To gain insights into the origin of the unstable flows in the system, we study the flow-induced sample turbidity using in-situ transmission-mode imaging [Fig. S4]. In WLM systems, sample turbidity occurs due to shear-induced local density variation over a range of length scales, enabling visible light scattering[chen2004flow, fardin2009taylor, ganapathy2008spatiotemporal2]. We perform these experiments in the separate motor transducer (SMT) mode where the pin is held fixed and the bottom sample cell is rotated at a desired angular velocity with the help of the bottom motor. Figs. 2(b)-(d) show the snapshots of the sample turbidity in steady-state corresponding to fixed applied velocities of 2.2 mm/s, 9 mm/s, and 15 mm/s, respectively. For v = 2.2 mm/s, we do not see any flow-induced structure, and the image shows a uniform darkness. For v = 9 mm/s, we observe the formation of a long, bright, extended wake that remains attached to the pin. The bright band-like structure remains stable under flow without showing any significant length variation. However, at the tail end of the stable wake, the presence of some local dynamics may result in weak temporal undulations in force [Fig. S13]. Nonetheless, our turbidity measurements are not sensitive enough to capture such dynamics of the wake. Interestingly, right after the beginning of the plateau region of the F-v plot, such a bright band starts to show continuous attachment-detachment events with the pin [Fig. 2(d), see movie 2]. The onset velocity of the pin to observe such events correlates with that of the stress fluctuation events in a one-to-one manner. This indicates that the attachment/detachment dynamics of the flow-induced structures are intricately related to the dynamics of steady-state force fluctuations. Interestingly, such sample turbidity under shear (Fig. 2b - 2d) does not depend on the polarization of light. We have tested this by measuring turbidity by placing a linear polarizer in front of the light source and an analyzer in front of the camera. We observe that the sample turbidity for a particular velocity of the pin (in the force-plateau region) remains unchanged for different combinations of the optical axes of the polarizer and the analyzer. Our observations are consistent with the sample turbidity obtained from bulk rheology measurements on the same sample, where no polarization dependence is observed [ghadai2023origin].
@@
Next, to establish a correlation between the attachment detachment dynamics with the observed force fluctuations, we map out the velocity distribution in the system using the particle imaging velocimetry (PIV) technique. As the micellar solution is transparent, 1.5 wt% of polystyrene tracer particles of size varying between 20-50 are added to the system to generate the contrast required for PIV. We confirm that adding such a small amount of tracer particles does not modify the rheological properties of the WLM system under study [Fig. S1]. For visualization, we image the top surface of the sample with the pin held fixed and the bottom geometry rotated [Fig. S5]. The recorded images are analyzed using the PIV Lab algorithm using MATLAB software[Thielicke_2021]. Fig. 3(a) shows the typical snapshot of the raw image of the sample-air interface seeded with tracer particles. Here we note that due to a difference between the radial distance of the two walls of the annular geometry from the axis of rotation, even without the presence of the pin the azimuthal velocity increases linearly from the inner to the outer wall. Fig. 3(b) shows the spatial distribution of velocity in the steady state for v = 2.5 mm/s. The velocity distribution in the presence of the pin (depicted as a black circle in the velocity distribution map) is almost similar to that without the pin. This indicates that for low velocities well below the onset of the plateau of the vs curve, the pin perturbs the system only locally, as also seen by a small low-velocity region around the pin [Fig. 3(b)]. When the applied velocity is increased to 12 mm/s (just below the onset of the plateau shown in Fig. 3(c)), a wake having a long, tail-like stationary structure is formed that remains connected to the probe and has much smaller velocity magnitude compared to the other portion of the sample. Interestingly, in the case of Newtonian fluids like glycerol, we observe that the presence of the pin perturbs the velocity distribution of the system only locally even if the velocity is varied over a wide range [Fig. S18]. This marks the striking difference between the flow fields around the pin for WLM system as compared to a high-viscosity Newtonian liquid. As mentioned earlier, we obtain large temporal fluctuations of F for an applied velocity that lies in the early plateau region of the curve where the features of the fluctuations are saw-tooth type. In Fig. 3(d), we show a zoomed-in version of a typical fluctuation event for v = 15 mm/s showing a slow build-up followed by a sharp drop in force. As indicated in the figure, we mark different regions in the time series as i, ii, iii, and iv. The corresponding spatial distributions of velocity are shown in Fig. 3 (panels i - iv). Just before the sharp force drop (region i), the spatial distribution of velocity shows the formation of a flow-induced long wake-like structure of a low-velocity band behind the pin (similar to the stable wake obtained for v = 12 mm/s). Such a stable wake quickly detaches from the pin and starts to move away from the pin with a much higher velocity compared to the velocity of the moving sample cell. At this point, force also starts to drop. Subsequently, the detached wake disturbs the sample around and pushes until a maximum speed of the sample is reached (iii and iv) and also reaches a minimum value. After this, the formation of a stable wake connected with the pin again takes place. The gradual increase in takes place correlating with the gradual elongation of the wake-like structure until the next force-drop event sets in. Such events continue as long as the sample cell velocity is maintained. Our results establish that the force fluctuation events and the formation and detachment of flow-induced structures are correlated in a one-to-one manner.
@@
In the increasing branch of the saw-tooth waveform for [Fig. 4(a)], we find that the length of the wake attached to the pin also increases with time till the maximum value of the force is reached (before the sharp force drop). Here, we define the length of the wake as the length of the region along the direction of azimuthal velocity over which the flow field gets altered due to the presence of the pin. Physically, this length scale signifies the region of influence of the pin for a given driving velocity and is obtained from the PIV measurement [panel i and ii in Fig. 4(a)]. We denote the maximum length of the wake formed within a saw-tooth waveform as . We further define a rise-time as the time required for to rise monotonically from a local minimum value to reach a maximum before it starts to drop abruptly as shown in Fig. 4(a). Earlier studies indicate that changing the degree of screening by varying the salt concentrations can tune the flow-concentration coupling in CTAT WLM[ganapathy2006intermittency]. We observe that a small variation of salt concentration (30 mM to 100 mM NaCl) does not change the nature of the curve except for slight shifting of the plateau onset [Fig. S10]. We also find a similar saw-tooth wave pattern close to the plateau onset for different salt concentrations. We construct a length scale from the force rise time: , where is the set velocity of the annular geometry at the location of the pin. We observe that (obtained from PIV measurement) and (obtained from the saw tooth behaviour of time-varying force) are very close to each other and vary in a correlated fashion as shown in Fig. 4(b) for a range of salt concentrations. Interestingly, we observe from Fig. 4(b) that the length of the wake before detachment increases with increasing salt concentration. Our results highlight that the rise of the force on the probe originates from the stretching of the micellar structure in the wake and the detachment of the structure from the pin results in the sharp drop in the force. Although close to the plateau region of the curve, the nature of most fluctuation events is quite close to the saw-tooth shape, few of the events are more complex (e.g. the first event shown in Fig. 4(a)) showing local non-monotonicity. In such cases, the wake region shows a partial detachment. We ignore such events in our correlation plot in Fig. 4(b). Interestingly, we observe that such partial detachment events become progressively more frequent as we enter deeper inside the plateau region. This observation rationalizes the complex waveform observed in the deep plateau region giving rise to more turbulent/chaotic flows. Apart from the change in salt concentration, temperature also plays a significant role in controlling the structural and mechanical properties of WLM. We find that the overall nature of the curve and the nature of fluctuations remain similar over a range of temperatures from to that we probe. We observe that for our system, the plateau onset shifts towards higher velocity with increasing temperature [Fig. S9]. However, estimating the length of the wake using PIV analysis at different temperatures remains outside the scope of the present study due to the technical limitation of controlling the sample temperature in twin-drive mode.
@@
The extensional deformation response of the system at room temperature of is probed using a custom made dripping on to substrate (DoS) rheometry technique following the standard protocol mentioned in the literature[dinic2015extensional, dinic2017pinch, omidvar2019detecting]. We observe the time evolution of a liquid bridge between a modified substrate and the needle that pinches off due to the role of strong capillary force. Fig. 6(a) shows the snapshots of the liquid bridge at different times for 2wt% CTAT + 100 mM based wormlike micellar system at room temperature of . We record the time evolution of the minimum neck diameter() of the unstable liquid bridge normalized w.r.t. the needle diameter () and extract the extensional relaxation time by fitting an exponential curve () as predicted by the elasto-capillary model[entov1997effect]. Fig. 6(b) shows the temporal evolution of the normalized mid-filament diameter for 2wt% CTAT system with variation of NaCl concentrations (see Fig. S21 for snapshots in the case of 30 mM and 50 mM NaCl) and the dashed lines are the fits using elasto-capillary model. The data is reproducible over multiple set of experiments [Fig. S22]. The extensional relaxation time obtained from DoS fittings over 4 repeated measurements is plotted w.r.t. the variation of NaCl concentrations (see inset of Fig.6(b)).
@@
Interestingly, 2 wt % CTAT + 100 mM NaCl does not show shear banding (before unstable/time-varying flows, the velocity profile remains linear). Such a non-banding nature of flow is also expected from the significant slope of stress plateau ( 0.3) in the flow curve and is also confirmed by direct imaging and PIV analysis [ghadai2023origin]. At high , the system directly enters into an unstable flow regime with random time-varying changes in the flow profile. However, in our pin-driving set-up, the force-velocity curve shows a very flat plateau, indicating the possibility of shear-banding. Indeed, our PIV measurement clearly shows that significant non-linearity in the flow profile develops near the plateau onset at much lower driving velocities beyond which the time-varying flow develops. Such a non-linear flow profile originates from the stable wake formation. The estimated value of Reynolds number (, where is the density, is the driving velocity, is the pin diameter, is the zero shear viscosity) at the plateau onset is much smaller than 1 () signifying that the elasticity plays a major role. The extensile Deborah number [moss2010flow, mohammadigoushki2016sedimentation] (, where is the maximum extension rate obtained from PIV measurement) at the plateau onset is 1. We further note that the shear ( ) and extensional () Weissenberg numbers ( and respectively denote the maximum shear and extension rates close to the pin) at the onset of time-varying flow in pin-driving set-up are also close to unity (, = 3). This onset observed for the pin driving set up is significantly lower than the onset of unstable flow in the Taylor-Couette geometry for the same system () [ghadai2023origin]. Nonetheless, in both cases, the instability is driven by fluid elasticity as . We note that alone may not be sufficient to characterize the onset of instabilities, as depends on flow geometries and streamline curvature. We also perform a Pakdel-McKinley analysis for our system to compare the onset of elastic instability [mckinley1996rheological, pakdel1996elastic] (see supplementary information for the calculations). We observe that for Taylor-Couette geometry, such analysis yields an onset value very close to our experimental observation, considering an upper-convected Maxwell model for the system. However, for the pin driving setup, the relevant geometry is the flow past a cylinder. Such an analysis is considerably more complicated, as one needs to estimate additional variables either from numerical modeling or by changing the parameters of the flow geometry. Furthermore, the baseline curvature of the streamlines that come from the circular shape of the sample cell can add further complexities. Exploring these directions remains outside the scope of the present study, but constitutes an important future direction. We believe that the main reason behind the difference in the nature of the flow curves obtained in pin driving and Taylor-Couette set up stems from the fact that in Taylor-Couette geometry, the flow is a simple shear flow to a good approximation. On the other hand, in pin driving, the flow has both shear and extensional components. The coupling between these different modes of deformations through the complex micellar structure can also play a role.
@@ IV Conclusions
In summary, we investigate the motion of a probe driven through a wormlike micellar fluid at a controlled velocity using a home-built shear cell coupled with in-situ flow visualization. For the first time, we correlate temporal stress fluctuations on the probe and the local shear and extensional deformations in a one-to-one fashion. We find that the flat force plateau obtained in the present setup differs from the finite slope of the stress plateau obtained in bulk rheology measurements for the same WLM system. We find a distinct saw-tooth feature of fluctuations at the onset of the instabilities. Similar features are not observed during bulk rheological measurements of the same system. We also observe some differences in the nature of the flow curve and the onset for stress fluctuations in the pin-driving set up as compared to those obtained for a Taylor-Couette geometry. Such differences can arise from the fact that both shear and extensional deformations play an important role in pin-driving as opposed to the simple shear flow in standard rheological geometries like Taylor-Couette. Interestingly, the context of flow past a cylinder is geometrically quite similar to the present setup, with an additional complexity of curvature of the sample cell in the pin-driving geometry. There have been detailed studies of flow instability for both Newtonian, as well as, viscoelastic shear-thinning and Boger fluids in the context of flow past a cylinder [mckinley1993wake, sheridan1997flow]. For viscoelastic flows, elastic wake instability and the formation of 3-D cellular structure along the length of the cylinder have been observed. Our set up can also be implemented in these cases to not only incorporate flow visualization but also to measure the mechanical force/stress reliably with good resolutions. However, in the current scenario, the pin completes one full rotation in less than 1 s for the pin velocity greater than 450 mm/s. For our WLM system, this sets the upper limit of the velocity as the shear relaxation time is 1 s. If the pin is driven with a higher velocity, the pin retraces its path before the sample fully relaxes. The force response will be dominated by the memory effect in the system, which hinders reliable mechanical measurements at such high driving velocities. Even at such high velocities, the Reynolds number just crosses 1 for Newtonian fluid glycerol. Thus, we need to go to a much higher velocity to induce inertia-driven instability in the case of glycerol. The reliability of force measurement and the high frame rate requirement need further optimization of the set up. Nonetheless, these directions will be very interesting to explore in the context of the motion of small probes in different viscoelastic systems.
Our study reveals the formation of flow-induced stable wakes just before the onset of force fluctuations. At higher driving velocities, the attachment detachment dynamics of the shear-induced micellar structures in the wake region with the probe remains strongly correlated with the stress fluctuations in the system. In the wake region, we get flow-induced turbidity which is visible even under the ambient light. This turbid band forms due to the scattering of light from the shear-induced structures in the non-linear flow regime. Interestingly, the scattered intensity is not sensitive to the polarization of light which we verify using different orientations of optical axes of a linear polarizer-analyzer pair. Understanding the microscopic origin of the sample turbidity will be an interesting future research direction. Careful measurement of sample birefringence might give us important insights. In the present study, we probe the stress fluctuations under a controlled velocity. Our preliminary data on controlled stress measurements reproduce the damped harmonic oscillations and bursts observed in the ball drop experiments in the CTAT-NaCl systems[kumar2012oscillatory]. Hence, stress/force-controlled measurements can yield significantly different temporal features of fluctuations in comparison to rate/velocity-controlled measurements. Our experiments can probe burst statistics much better as the conventional ball drop experiments are limited by the finite tube length. However, we must remember that the probe is held rigidly in the current setup which does not allow any rotational or lateral motion of the probe. However, in the case of ball-drop experiments, no such restrictions are present. This can affect the dynamics. Introducing such flexibility in our set-up will be an interesting direction for future studies.
@@
See the supplementary material for information about different imaging set-ups, data analysis and additional measurements.